### v2

# PRIMEdit: Probability Redistribution for Instance-aware Multi-object Video Editing with Benchmark Dataset

Abstract
Recent AI-based video editing has enabled users to edit videos through simple text prompts, significantly simplifying the editing process. However, recent zero-shot video editing techniques primarily focus on global or single-object edits, which can lead to unintended changes in other parts of the video. When multiple objects require localized edits, existing methods face challenges, such as unfaithful editing, editing leakage, and lack of suitable evaluation datasets and metrics. To overcome these limitations, we propose Probability Redistribution for Instance-aware Multi-object Video Editing (PRIMEdit). PRIMEdit is a zero-shot framework that introduces two key modules: (i) Instance-centric Probability Redistribution (IPR) to ensure precise localization and faithful editing and (ii) Disentangled Multi-instance Sampling (DMS) to prevent editing leakage. Additionally, we present our new MIVE Dataset for video editing featuring diverse video scenarios, and introduce the Cross-Instance Accuracy (CIA) Score to evaluate editing leakage in multi-instance video editing tasks. Our extensive qualitative, quantitative, and user study evaluations demonstrate that PRIMEdit significantly outperforms recent state-of-the-art methods in terms of editing faithfulness, accuracy, and leakage prevention, setting a new benchmark for multi-instance video editing.
Figure 1: Given a video, instance masks, and target instance captions, our PRIMEdit framework enables faithful and disentangled edits in (a) single- and (b)-(c) multi-instance levels, as well as an applicability to more fine-grained (d) partial instance level without the need for additional training. Unlike previous methods, our PRIMEdit does not rely on global edit captions, but leverages individual instance captions. Each object mask is color-coded to match its corresponding edit caption. Zoom-in for better visualization.
1 Introduction
The popularity of short-form videos on social media has grown significantly [53, 31]. However, editing these videos is often time-consuming [50] and can require professional assistance [37]. These challenges have driven the development of AI-based video editing (VE) tools [31], supported by advances in generative [46, 68, 4] and visual language models [43]. These tools allow users to specify desired edits with simple text prompts [2, 42], making the editing process faster and more accessible.
Recent VE methods often leverage pre-trained text-to-image (T2I) models [46]. Compared to alternative approaches, such as training models on large-scale datasets [73, 12] or fine-tuning on single videos [59, 36, 2], zero-shot methods [72, 14, 30, 63, 20, 42, 28, 8] continue to gain traction due to their efficiency and the availability of pre-trained T2I models. Most zero-shot approaches focus on global editing that modifies the entire scene [72, 63] or single-object editing that can unintentionally affect other parts of video [7, 36]. In some cases, however, users may require precise editing of particular objects without altering other parts of the video, such as replacing explicit content (e.g., cigarettes) to create family-friendly versions [70].
Figure 2: Limitations of previous SOTA methods. (a) ControlVideo [72] relies on single global captions, and (b) GAV [28] depends on bounding box conditions that can sometimes overlap. Both are susceptible to unfaithful editing (red arrow) and attention leakage (blue arrow).
Local VE aims to address this problem by accurately manipulating specific objects in videos. However, adapting pre-trained T2I models for this task is challenging since they lack fine-grained control and require additional training [73] or attention manipulation [65, 36] to enable spatial control. The problem is further exacerbated when the models need to perform multiple local edits simultaneously using a single long caption as observed in Fig. 2-(a) for ControlVideo [72]. This approach often leads to: (i) unfaithful editing (e.g., the red bag not transforming into a yellow purse) and (ii) attention leakage [62] where the intended edit for a specific object unintentionally affects other object regions (e.g., both wall and statue becoming yellow).
Recently, Ground-A-Video (GAV) [28] demonstrated that simultaneous multi-object VE is feasible using a T2I model [46] fine-tuned on grounding conditions [33]. However, GAV suffers from attention leakage [62], especially when the objects’ bounding boxes overlap (e.g., the statue also gaining flower crown in Fig. 2-(b)). VideoGrain [65] attempts to address the leakage by instead using object masks and attention modulation [32]. However, VideoGrain’s use of attention modulation in the spatio-temporal attention (STA) imposes significant memory and computation bottlenecks, particularly as the number of video frames increases see Supplemental). VideoGrain also relies on global edit captions to describe the edited scenes, which may be unavailable or challenging to generate in scenes with a large number of objects.
Furthermore, both GAV [28] and VideoGrain [65] have been tested on limited datasets, insufficient for comprehensive testing of multi-instance VE across diverse viewpoints, instance sizes, and numbers of instances. The global editing metrics used in these methods further fail to precisely measure local editing quality, essential for multi-instance VE.
From the above, four critical challenges persist in multi-instance VE: (i) Attention leakage. The absence of local control in pre-trained T2I models, the imprecise input conditions [28] and the use of a single global edit caption [72] hinder previous methods from effectively disentangling edit prompts; (ii) Unfaithful editing. The lack of techniques to enhance editing faithfulness in multi-instance VE tasks often results in inaccurate editing; (iii) Lack of fast and intuitive multi-instance VE methods. While modulation in [65] enables per-object editing, it slows down the generative editing process. Moreover, its reliance on full-scene descriptions makes editing cumbersome, especially in highly cluttered videos. (iv) Lack of evaluation dataset and metrics. Recent methods [65, 28] are tested on limited datasets using metrics inadequate for evaluating local VE quality.
To overcome these, we present our Probability Redistribution for Instance-aware Multi-object Video Editing framework (PRIMEdit), a more intuitive zero-shot multi-instance VE method to achieve faithful edits and reduce attention leakage by disentangling multi-instance edits. Our PRIMEdit can easily adopt existing T2I and T2V models into a unified framework and achieve multi-instance editing capabilities through two key modules: (i) To enhance editing faithfulness and increase the likelihood of objects appearing within their masks, we introduce Instance-centric Probability Redistribution (IPR, Sec. 3.2) in the cross-attention layers; and (ii) Inspired by prior works [55, 48], we design Disentangled Multi-instance Sampling (DMS, Sec. 3.3) which can significantly reduce attention leakage. Moreover, we evaluate PRIMEdit on our newly proposed benchmark dataset, called MIVE Dataset, of 200 videos using standard metrics and a novel metric, the Cross-Instance Accuracy (CIA) Score, to quantify attention leakage. Our contributions are as follows:
  • We propose a novel zero-shot multi-instance video editing framework, called PRIMEdit, which enables fast and intuitive multi-instance editing for videos;
  • We propose to disentangle multi-instance video editing through our (i) Instance-centric Probability Redistribution (IPR) that enhances editing localization and faithfulness and (ii) Disentangled Multi-instance Sampling (DMS) that reduces editing leakage;
  • We propose a new evaluation benchmark that includes a novel evaluation metric, called the Cross-Instance Accuracy (CIA) Score, and a new dataset, called MIVE Dataset, consisting of 200 videos with varying numbers and sizes of instances accompanied with instance-level captions and masks. The CIA Score is designed to quantify attention leakage in multi-instance editing tasks;
  • Our extensive experiments validate the effectiveness of our PRIMEdit in disentangling multiple edits across various instances and achieving faithful editing with better temporal consistency, significantly outperforming the most recent SOTA video editing methods.
Figure 3: The overall framework of our PRIMEdit, given $M$ number of multi-instance captions $c_{i}$ with corresponding instance masks $\boldsymbol{m_{i}}$ for editing. Our Disentangled Multi-instance Sampling (DMS, Sec. 3.3) consists of series noise sampling (SNS, yellow box), latent fusion (green box) to fuse different instance latents, re-inversion (purple box) to harmonize the latents after fusion, and parallel noise sampling (PNS, blue box). In addition, our Instance-centric Probability Redistribution (IPR, Sec. 3.2) provides better spatial control.
2 Related Works
Zero-shot text-guided video editing. Recent advancements in diffusion models [25, 51, 52] have accelerated the evolution of both text-to-image (T2I) [44, 40, 47, 15, 46] and text-to-video (T2V) [4, 69, 57, 26, 22, 18, 54] models for generative tasks. These breakthroughs led to the development of numerous video editing (VE) frameworks, where pre-trained models serve as backbones. Most VE methods [42, 59, 20, 14, 30, 28, 72] rely on pre-trained T2I models [46, 33, 71], as early T2V models were either not publicly accessible [5] or computationally expensive to run [54]. To facilitate VE, recent methods have either fine-tuned T2I models on large video datasets [73, 56, 18], optimized on a single input video [59, 36], or leveraged zero-shot inference [72, 14, 30, 63, 20, 28, 42, 65]. Our work falls within zero-shot VE, allowing editing without additional training.
Local video editing based on image generation and editing. Similar to image generation and editing, the need for fine-grained control in videos has led to the development of several local VE methods [73, 36, 28, 65]. These methods usually adopt spatial control techniques from image generation such as training additional adapters [71, 1, 33, 66, 55] or employing zero-shot techniques, e.g., attention manipulation [32, 23], optimization [3, 9, 10, 61], and multi-branch sampling [48]. In particular, AVID [73] uses masks and retrains the Stable Diffusion (SD) inpainting model [46], while Video-P2P [36] leverages an attention control method [23], enabling both to achieve finer control.
Extending the local control for multi-instance VE scenarios is relatively underexplored, with few works [28, 65] tackling this challenge by similarly adopting image generation techniques [33, 32] to localize VE. GAV [28] leverages GLIGEN [33] and integrates gated self-attention layers within the transformer blocks, allowing for spatial control. However, GLIGEN requires retraining the gated self-attention layers for every new T2I model (e.g. SDv1.5 and SDv2.1), reducing its flexibility. VideoGrain [65] uses masks and attention modulation [32] for the spatio-temporal attention (STA) and cross-attention layers. However, manipulating the STA increases computational costs as the number of edited frames increases. Our work introduces a novel redistribution technique that manipulates values only within the cross-attention layers, allowing for seamless integration with any T2I or T2V model and spatial control with minimal computational overhead.
Text-to-video models. A growing trend in T2V generation is training a Diffusion Transformer (DiT) [5, 11] on largescale video datasets. Prior to DiTs, T2V models were built by adding temporal layers to pre-trained T2I models [4, 19, 21] and fine-tuning them on video datasets for temporal modeling as it is more efficient.
Recently, several works [73, 67] have started adopting fine-tuned T2V models in VE to achieve better temporal consistency. AVID [73] extends a pre-trained T2I model [46] to T2V by inserting motion modules and finetunes on a large video dataset to enable fine-grained video editing. DMT [67] introduced a space-time feature loss derived from a T2V model [54, 6] to transfer motion from objects in a reference video to target edit objects in a zero-shot manner. Similarly, we use a pre-trained publicly available T2V model [21] to achieve better temporal consistency.
3 Proposed Method
3.1 Overall Framework
In this work, we tackle multi-instance video editing (VE) by disentangling the edits for multiple instances. Given a set of $N$ input frames $\boldsymbol{f}=f^{1:N}$ and a set of $M$ instance target edits $\boldsymbol{g}=\{g_{i}\}_{i=1}^{M}$, where each target edit $g_{i}=\{\boldsymbol{m_{i}},c_{i}\}$ consists of instance masks $\boldsymbol{m_{i}}=m_{i}^{1:N}$ and corresponding edit captions $c_{i}$, we modify each instance $i$ based on $c_{i}$, ensuring that the region outside the masks $m_{i}^{1:N}$ remain unedited.
Fig. 3 illustrates the overall framework of our PRIMEdit. PRIMEdit falls under the category of inversion-based VE [42] (see Supp. for preliminaries). Starting with the initial latents $z_{0}^{1:N}=\mathcal{E}(f^{1:N})$ generated using the VAE encoder $\mathcal{E}$ [46], we employ the 3D U-Net [46, 21] to invert the latents using DDIM Inversion [51]. This yields a sequence of inverted latents, $\{\boldsymbol{\tilde{z}_{t}}\}_{t=0}^{T}=\{\tilde{z}_{t}^{1:N}\}_{t=0}^{T}$, that we store for later use with $T$ representing the number of denoising steps.
The vanilla cross-attention within the U-Net [46] cannot ensure the target edit to stay within $\boldsymbol{m_{i}}$ (see Fig. 6-(b)). To address this, we introduce Instance-centric Probability Redistribution (IPR, Sec. 3.2) in the cross-attention, boosting the accuracy of edit placement inside $\boldsymbol{m_{i}}$.
For the sampling, we introduce the Disentangled Multi-instance Sampling (DMS, Sec. 3.3), inspired by image generation methods [55, 48], to disentangle the multi-instance VE process, and minimize attention leakage. Each instance is independently modified using Series Noise Sampling (SNS, blue box in Fig. 3), and the multiple denoised instance latents are harmonized through Parallel Noise Sampling (PNS) preceded by Latent Fusion and Re-inversion (green, yellow, and red boxes, respectively, in Fig. 3). We employ ControlNet [71] during sampling, with depth maps $\boldsymbol{d}=d^{1:N}_{i}$ obtained via MiDas [45] as conditions.
3.2 Instance-centric Probability Redistribution
Figure 4: A comparative illustration of our IPR versus others (top) and details of our IPR (bottom).
Our sampling requires the edited objects to appear within their masks, highlighting the need for spatial control in the U-Net because the vanilla cross-attention [46] struggles to localize edits (see Fig. 6-(b)). To address this, we propose Instance-centric Probability Redistribution (IPR), inspired by attention modulation [32]. Our IPR enables faithful editing while operating exclusively within the cross-attention layers, avoiding the significant computational overhead associated with using [32] in both the spatio-temporal and cross-attention layers [65] to produce high-quality results (see Fig. 6-(c)). Fig. 4 shows our IPR (bottom part) and a comparative illustration (top part) versus others.
In the bottom part of Fig. 4, we focus on single instance editing with a target caption $c=c_{i}$. The caption $c$, with $n$ text tokens, is encoded into text embeddings using a pretrained CLIP model [43], which are then used as the key $\boldsymbol{K}$ in the cross-attention. Each value of key $K_{j=\{S,\boldsymbol{T},E,\boldsymbol{P}\}}$ corresponds to either a start of sequence $S$, multiple text $\boldsymbol{T}$, an end of sequence $E$, and multiple padding $\boldsymbol{P}$ tokens. The cross-attention map $\boldsymbol{A}$ between the query image features $\boldsymbol{Q}\in\mathbb{R}^{hw\times d}$ and $\boldsymbol{K}$ can be expressed as:
$\boldsymbol{A}=\text{Softmax}(\boldsymbol{Q}\boldsymbol{K}^{T}\mathbin{/}\sqrt{d})\in[0,1]^{hw\times n},$ (1)
[…]
In our IPR, we split $A_{i,j}$ of an instance into two sets depending on the mask $m$: outside the instance $\boldsymbol{A}_{\boldsymbol{O},j}=\{A_{i,j};m_{i}=0\}$ and inside the instance $\boldsymbol{A}_{\boldsymbol{I},j}=\{A_{i,j};m_{i}=1\}$. Our IPR is based on several experimental observations (see Supp. for details): (i) Manipulating the attention probabilities of the padding tokens $A_{i,j\in P}$ may lead to artifacts, so we avoid altering these; (ii) Increasing $S$ token’s probability decreases editing faithfulness, while decreasing it and reallocating those values to the $\boldsymbol{T}$ and $E$ tokens increases editing faithfulness. Thus, for $\boldsymbol{A}_{\boldsymbol{O},j}$, we zero out $\boldsymbol{T}$ and $E$ tokens’ probabilities, reallocating them to $S$ to prevent edits outside the mask (blue dashed box in Fig. 4, bottom). In addtition, we redistribute the attention probability for the object region inside the mask by reducing the value of $S$ and redistributing it to $\boldsymbol{T}$ and $E$ (red dashed box in Fig. 4, bottom). We reduce $\boldsymbol{A}_{\boldsymbol{I},j=S}$ by $\lambda_{S}$, which decays linearly to 0 from $t=T$ to $t=1$, formulated as:
[…]
where $W$ is a warm-up value that is linearly decayed from $\lambda$ to 0 in the early sampling steps $t<0.1T$ to increase the editing fidelity especially for small and difficult objects. Increasing $\lambda$ tends to enhance faithfulness but may introduce artifacts. We empirically found that $\lambda=0.5$ is the best trade-off between faithfulness and artifact-less results. Then, we redistribute $\lambda_{S}$ to $\boldsymbol{T}$ and $E$ with the ratios of $\lambda_{r}$ and $1-\lambda_{r}$, respectively. The attention probabilities, $\boldsymbol{A}_{\boldsymbol{I},j\in\boldsymbol{T}}$ and $\boldsymbol{A}_{\boldsymbol{I},j=E}$, of $\boldsymbol{T}$ and $E$ are updated as follows:
$\begin{gathered}\boldsymbol{A}_{\boldsymbol{I},j\in\boldsymbol{T}}=\boldsymbol{A}_{\boldsymbol{I},j\in\boldsymbol{T}}+\lambda_{S}\cdot\lambda_{r}\mathbin{/}N_{\boldsymbol{T}}\\ \boldsymbol{A}_{\boldsymbol{I},j=E}=\boldsymbol{A}_{\boldsymbol{I},j=E}+\lambda_{S}\cdot(1-\lambda_{r})\end{gathered}$ (3)
where $N_{\boldsymbol{T}}$ denotes the number of text tokens. Increasing $\lambda_{r}$ may lead to enhancing the editing details for certain tokens, while decreasing it may increase the overall editing fidelity. $\lambda_{r}=0.5$ is empirically set for all experiments.
3.3 Disentangled Multi-instance Sampling
To disentangle the multi-instance VE process and reduce attention leakage, we propose our Disentangled Multi-instance Sampling (DMS). As shown in Fig. 3, DMS comprises of two sampling strategies: (1) Series Noise Sampling (SNS) shown in yellow box and (2) Parallel Noise Sampling (PNS) shown in blue box preceded by latent fusion (green box) and re-inversion (purple box). In the SNS, we independently edit each instance $i$ using its target caption $c_{i}$ and its mask $\boldsymbol{m_{i}}$ by first estimating each instance noise $\hat{n}_{i}^{1:N}$ as:
$\hat{n}_{i}^{1:N}=\epsilon_{\theta}(\{z_{t}^{1:N}\}_{i},c_{i},\boldsymbol{m_{i}},e,t),$ (4)
and the background noise $\tilde{n}^{1:N}$ using the inverted latents as:
$\tilde{n}^{1:N}=\epsilon_{\theta}(\tilde{z}_{t}^{1:N},c_{\varnothing},t).$ (5)
Then, we fuse the instance noise $\hat{n}_{i}^{1:N}$ with the inverted latent noise $\tilde{n}^{1:N}$ to obtain the edited instance noise $\{\bar{n}_{t-1}^{1:N}\}_{i}$ using masking as:
$\{\bar{n}_{t-1}^{1:N}\}_{i}=\hat{n}_{i}^{1:N}\cdot\boldsymbol{m_{i}}+\tilde{n}^{1:N}\cdot(1-\boldsymbol{m_{i}}).$ (6)
We then perform one DDIM [51] denoising step from $t\rightarrow t-1$ to obtain the instance latents $\{\hat{z}_{t-1}^{1:N}\}_{i}$ as:
$\{\hat{z}_{t-1}^{1:N}\}_{i}=\text{DDIM}(\{\hat{z}_{t}^{1:N}\}_{i},\{\bar{n}_{t-1}^{1:N}\}_{i},t).$ (7)
SNS requires the edited instance $i$ to appear within $\boldsymbol{m_{i}}$, as the background is replaced with the inverted latent noise at each step to ensure it remains unchanged. Our proposed IPR (Sec. 3.2) provides this necessary spatial control.
While we can perform multi-instance VE by solely using SNS until the end of denoising, the edited frames suffer from temporal inconsistency (see Tab. 3). Thus, we propose two approaches to mitigate this:
(i) we fuse the multiple instance latents $\{\dot{z}_{t}^{1:N}\}$ sampled independently from SNS as follows:
$\dot{z}_{t}^{1:N}=\Sigma_{i=1}^{M}(\{\hat{z}_{t}^{1:N}\}_{i}\cdot\boldsymbol{m_{i}})+\tilde{z}_{t}^{1:N}\cdot\boldsymbol{m_{B}},$ (8)
where $\boldsymbol{m_{B}}=\boldsymbol{1}-\Sigma_{i=1}^{M}\boldsymbol{m_{i}}$ denotes a background mask;
(ii) we propose re-inversion using DDIM for $l$ steps after the latent fusion to obtain $\ddot{z}_{t+l}^{1:N}$ followed by PNS from timestep $t+l\rightarrow t$ to yield the final fused latent $\ddot{z}_{t}^{1:N}$.
We then perform PNS for the remaining denoising steps. The goal of our PNS is to harmonize the independent instance latents obtained from the SNS while still decoupling the sampling using the individual target captions. In this sampling strategy, we use the re-inverted fused latents $\ddot{z}_{t}^{1:N}$ to estimate each instance noise $\hat{n}_{i}^{1:N}=\epsilon_{\theta}(\ddot{z}_{t}^{1:N},c_{i},\boldsymbol{m_{i}},e,t)$ using the instance caption $c_{i}$. We still estimate the noise $\tilde{n}^{1:N}=\epsilon_{\theta}(\tilde{z}_{t}^{1:N},c_{\varnothing},t)$ for the background using the inverted latents $\tilde{z}_{t}^{1:N}$ and an empty caption $c_{\varnothing}$. We then combine $\hat{n}_{i}^{1:N}$ and $\tilde{n}^{1:N}$ to a single noise through masking, and perform one DDIM step to obtain the latents $\hat{z}_{t-1}^{1:N}$ as:
$\hat{z}_{t-1}^{1:N}=\text{DDIM}(\hat{z}_{t}^{1:N},\Sigma_{i=1}^{M}(\hat{n}_{i}^{1:N}\cdot\boldsymbol{m_{i}})+\tilde{n}^{1:N}\cdot\boldsymbol{m_{B}},t).$ (9)
Finally, we decode $\hat{z}_{t-1}^{1:N}$ using the VAE decoder $\mathcal{D}$ to obtain the final edited frames $\hat{\boldsymbol{f}}=\mathcal{D}(\hat{z}_{0}^{1:N})$.
Dataset Name Number Number of Number of Number of Object Number of Instances Instance Instance Range of Average Instance
of Clips Frames per Clip Objects per Clip Classes per Object Class Captions Masks Mask Size Per Video (%)
TGVE [60] & TGVE+ [49] 76 $32-128$ 1-2 No Info 1-2 $\checkmark$ $\times$ No Masks
VideoGrain [65] 13 $16-32$ 1-5 7 1-2 $0.00\sim 92.34$
MIVE Dataset (Ours) 200 $12-46$ 3-12 110 1-20 $0.01\sim 98.68$
Table 1: Comparison between our multi-instance video editing dataset with other text-guided video editing datasets.
4 Evaluation Data and Metric
4.1 MIVE Dataset Construction
Existing datasets are unsuitable for multi-instance video editing (VE) tasks. TGVE [60] and TGVE+ [49] contain a limited number of videos, feature few objects with few instances per object class, and lack instance masks. Similarly, the dataset used in VideoGrain [65] includes only a few objects with limited instances per class. GAV’s [28] dataset is only partially available, and DAVIS [41] does not distinguish between multiple instances of the same object class, making it unsuitable for multi-instance VE.
To address this gap, we introduce the MIVE Dataset, a new evaluation dataset for multi-instance VE tasks. Our MIVE Dataset features 200 diverse videos from VIPSeg [39], a video panoptic segmentation dataset, with each video center-cropped to a $512\times 512$ region. Since VIPSeg lacks source captions, we use LLaVA [35] to generate captions and use Llama 3 [17] to summarize the captions to a fewer tokens. We then manually insert tags in the captions to establish object-to-mask correspondence. Finally, we use Llama 3 to generate the target edit captions by swapping or retexturing each instance similar to [73].
Table 1 compares our MIVE Dataset to other VE datasets. Compared to others, MIVE Dataset offers the most number of evaluation videos and the greatest diversity in terms of the number of objects per video and the number of instances per object class. Our dataset also demonstrates a wide range of instance sizes, from small objects covering as little as 25 pixels ($\sim$0.01%) per video to large objects occupying $\sim$98.68of the video. See Supp. for more details.
4.2 Cross-Instance Accuracy Score
SpaText [1] and InstanceDiffusion [55] assess local textual alignment using the Local Textual Faithfulness, the cosine similarity between CLIP [43] text embeddings of an instance caption and image embeddings of the cropped instance. While this measures text-instance alignment, it overlooks potential cross-instance information leakage, where a caption for an instance influences others. We observe that this score is sometimes higher for instances that should be unaffected by an instance caption (see Supp.).
To address this, we propose a new metric called Cross-Instance Accuracy (CIA) Score that we define as follows: For each cropped instance $i$, we compute the cosine similarity $S(I_{i},C_{j})$ between its CLIP image embeddings $I_{i}$ and the text embeddings $C_{j}$ for all $n$ instance captions $(i,j\in\{1,...,n\})$, producing an $n\times n$ similarity matrix given by $\mathbf{S}=[S(I_{i},C_{j})]$ for $i,j=1,2,...,n$. For each row in S, we set the highest similarity score to 1 and all others to 0, ideally resulting in a matrix with 1’s along the diagonal and 0’s elsewhere, indicating that each cropped instance aligns best with its caption. The CIA is calculated as the mean of diagonal elements as:
$\text{CIA}=(1\mathbin{/}n)\cdot\Sigma_{i=1}^{n}S(I_{i},C_{i}).$ (10)
5 Experiments
Method Venue Editing Local Scores Leakage Scores User Study
Scope LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ TC $\uparrow$ TF $\uparrow$ Leakage $\uparrow$
ControlVideo [72] ICLR’24 Global 0.9548 0.1960 0.4944 0.4963 0.5327 0.4242 0.0167 0.0656 0.0208
FLATTEN [14] ICLR’24 Global 0.9507 0.1881 0.2463 0.5109 0.8220 0.1314 0.3260 0.0594 0.1646
RAVE [30] CVPR’24 Global 0.9551 0.1869 0.3504 0.4950 0.7699 0.2188 0.0667 0.0354 0.0458
GAV [28] ICLR’24 Local, Multiple 0.9518 0.1895 0.3733 0.5516 0.8618 0.0914 0.0792 0.0938 0.1000
VideoGrain [65] ICLR’25 Local, Multiple 0.9423 0.2026 0.4798 0.5868 0.8929 0.0488 0.0802 0.1177 0.1542
PRIMEdit (Ours) - Local, Multiple 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586 0.3365 0.5687 0.4510
Table 2:  Quantitative comparison for multi-instance video editing. The best and second best scores are shown in red and blue, respectively.
Implementation details. We evaluate our PRIMEdit framework on our MIVE dataset (Sec. 4.1), editing 12-32 frames per video, and conduct our experiments on a single NVIDIA RTX A6000 GPU. We implement our PRIMEdit on top of a baseline that integrates Stable Diffusion [46] v1.5, AnimateDiff [21] motion modules and ControlNet [71] with depth [45] as our condition. We perform $T=50$ DDIM denoising steps after doing 100 inversion steps with an empty text following [14]. Our CFG [24] scale is 12.5. We apply our IPR in the first $10\%$ of the total denoising steps and switch to vanilla cross-attention [46] in the remaining steps. We perform SNS for the first $40\%$ of the denoising steps and perform PNS for the remaining steps. Our re-inversion steps $l$ is set to $l=2$.
Evaluation metrics. To evaluate the edited frames, we report standard metrics: (i) Global Temporal Consistency (GTC) [64, 30, 14, 63]: average cosine similarity of CLIP [43] image embeddings between consecutive frames, (ii) Global Textual Faithfulness (GTF) [64, 30, 14, 20]: average similarity between the frames and global edit prompts, and (iii) Frame Accuracy (FA) [64, 63]: percentage of frames with greater similarity to the target than the source prompt.
Global metrics assess overall frame quality but they overlook nuances in individual instance edits, essential for multi-instance tasks. To address this, we use Local Temporal Consistency (LTC), Local Textual Faithfulness (LTF) [55, 1], and Instance Accuracy (IA) computed using the cropped edited instances. Metric computation details are in the Supp. We also quantify cross-instance leakage through our proposed CIA (Sec. 4.2) and background leakage through SSIM [58] and LPIPS [13, 29] scores.
5.1 Experimental Results
Figure 5: Qualitative comparison for three videos (with increasing difficulty from left to right) in our MIVE dataset. (a) shows the color-coded masks overlaid on the input frames to match the corresponding instance captions. (b)-(d) use global target captions for editing. (e) uses global and instance target captions along with bounding boxes (omitted in (a) for better visualization). (f) uses masks and global and local target captions. Our PRIMEdit in (g) uses instance captions and masks. Unfaithful editing examples are shown in red arrow and attention leakage are shown in green arrow.
We compare our PRIMEdit against recent zero-shot video editing (VE) methods: (i) five global editing: ControlVideo [72], FLATTEN [14], RAVE [30], FreSCo [63], TokenFlow [20], and (ii) two multi-object editing: GAV [28] and VideoGrain [65]. We show results for the first 3 global and 2 multi-object video editing baselines in this main paper and present the rest in the Supp. We use (i) single global source and target captions for the global editing methods and (ii) a combination of global and local source and target captions along with bounding box and mask conditions for GAV and VideoGrain. Our PRIMEdit requires object masks and local target captions only. We exclude Video-P2P [36] since it edits one instance at a time, accumulating error when used in multi-instance editing scenarios (see Supp.). For each baseline above, we use its default settings.
Qualitative comparison. We present qualitative comparisons in Fig. 5. As shown, ControlVideo, GAV, and VideoGrain suffer from attention leakage (e.g. “yellow” affecting the “alien” in Video 1 of Fig. 5-(b), (e), and (f)) while FLATTEN, RAVE, and GAV exhibit unfaithful editing for all the examples (Fig. 5-(c) to Fig. 5-(e)). Moreover, ControlVideo exhibit a mismatch in editing instances as shown in Fig. 5-(b), Video 2 (incorrectly turning the man on the left to “woman in a red dress”). In contrast, our PRIMEdit confines edits within the instance masks, faithfully adheres to instance captions, and does not exhibit mismatch in instance editing.
Quantitative comparison. Table 2 shows the quantitative comparisons. We show only the local and leakage scores along with the user study results in this main paper and present the global scores in the Supp. The SOTA methods yield (i) low LTF and IA scores, indicating less accurate instance-level editing, and (ii) low CIA scores, reflecting significant attention leakage. In contrast, our PRIMEdit achieves the best scores in key multi-instance VE metrics (LTC, LTF, IA, and CIA) demonstrating its superior ability to maintain temporal consistency, fine-grained textual alignment, and instance-aware modifications. Note also that our PRIMEdit exhibits strong performance in background preservation metrics, achieving the best SSIM and second-best LPIPS scores without relying on an inpainting model like GAV [28]. This further underscores its effectiveness in isolating instance edits while preserving and preventing leakage in the background. We provide more qualitative comparisons in the demo videos in the Supp.
Methods LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
IPR No Modulation [46] 0.9564 0.2018 0.4920 0.6497 0.9007 0.0600
Dense Diffusion [32] 0.9530 0.2078 0.5845 0.6774 0.8977 0.0604
Ours, Full 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
DMS Only SNS 0.9503 0.2047 0.5288 0.6743 0.9063 0.0500
Only PNS 0.9542 0.2044 0.5311 0.6718 0.8993 0.0585
SNS + PNS (no re-inv) 0.9535 0.2051 0.5323 0.6728 0.9040 0.0525
Ours, Full 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
Table 3: Ablation study results on IPR (Sec. 3.2) and DMS (Sec. 3.3). SNS, PNS, and re-inv denotes Series Noise Sampling, Parallel Noise Sampling, and Re-Inversion, respectively. The best and second best scores are shown in red and blue, respectively.
User study. We conduct a user study to compare our PRIMEdit with the SOTA baselines, and report the results in Tab. 2. We select 30 videos from our dataset covering diverse scenarios (varying number of objects per clip, number of instances per object class, and instance size). We task 32 participants to select the method with the best temporal consistency (TC), textual faithfulness (TF), and minimal editing leakage (Leakage). Our PRIMEdit demonstrates clear superiority against previous SOTA methods with 33.65% chance of being selected as the best for TC, 56.87% for TF, and 45.10% for the least amount of leakage. Details of our user study are in the Supp.
5.2 Ablation Studies
In our ablation studies, we report key multi-instance VE metrics in Tab. 3 and provide the Global Scores in the Supp.
Figure 6: Ablation study on IPR (Sec. 3.2).
Ablation study on IPR. We conduct an ablation study on our IPR, with qualitative results presented in Fig. 6 and quantitative results in Tab. 3. As shown in Fig. 6-(b), the model tends to reconstruct the input frames when modulation is omitted. The reconstruction thereby preserves the high LTC of the input but suffers from unfaithful editing (low LTF and IA) and leakage (low CIA). Introducing cross-attention modulation via DenseDiffusion [32] reduces LTC but improves faithfulness and minimizes leakage. However, as presented in the green box in Fig. 6-(c) and [56], it struggles to generate visually appealing results when the computationally expensive spatio-temporal attention modulation is absent. In contrast, our IPR effectively reduces background artifacts (best SSIM and LPIPS), maintains temporal consistency (second-best), enhances edit faithfulness (second-best LTF and IA), and minimizes leakage (second-best CIA), all while operating solely on the cross-attention. We provide cross attention visualization in the Supp.
Figure 7: Ablation study on DMS (Sec. 3.3).
Ablation study on DMS. We also conduct an ablation study on our DMS, presenting qualitative results in Fig. 7 and quantitative results in Tab. 3. From Fig. 7 and Tab. 3, we observe the following: (i) using only SNS prevents excessive leakage (high CIA) but suffers from temporal inconsistency (lowest LTC) as shown in Fig. 7-(b), green arrow; (ii) using only PNS improves LTC as objects are fused early in the denoising process but introduces some leakage (decrease in CIA), as seen in Fig. 7-(c), red arrow, where the green color from the plants leaks onto the man’s face; (iii) performing SNS followed by PNS mitigates the issues observed in (i) and (ii), improving both CIA and LTC; and (iv) adding re-inversion improves IA and boosts LTC by harmonizing instances with a slight decline in CIA. Overall, our final model strikes a balanced trade-off between LTC, faithfulness, leakage, and background preservation.
6 Conclusion
In this work, we introduce PRIMEdit, a novel multi-instance video editing framework featuring Instance-centric Probability Redistribution (IPR) and Disentangled Multi-instance Sampling (DMS). Our method achieves faithful and disentangled edits with minimal attention leakage, outperforming existing SOTA methods in both qualitative and quantitative analyses on our newly proposed MIVE dataset. We further propose a new Cross-Instance Accuracy (CIA) Score to quantify attention leakage. Our user study supports PRIMEdit’s robustness and effectiveness with participants favoring our method.
Supplementary Material
A Preliminaries
We briefly introduce how inversion-based video editing is achieved in this subsection. Given a set of input frames $f^{1:N}$, each frame $f^{i}$ is encoded into a clean latent code $z_{0}^{i}=\mathcal{E}(f^{i})$ using the encoder $\mathcal{E}$ of the Latent Diffusion Model (LDM) [46]. DDIM inversion [51, 15] is applied to map the clean latent $z_{0}^{i}$ to a noised latent $\hat{z}_{T}^{i}$ through reversed diffusion timesteps $t:1\rightarrow T$ using the U-Net $\epsilon_{\theta}$ of the LDM:
[…]
where $\alpha_{t}$ denotes a noise scheduling parameter [46] and $I_{t}=(\hat{z}_{t},t,c,e)$ represents the noisy latent at timestep $t$ with text prompt $c$ as input. After inversion, the latents $\hat{z}_{t}^{i}$ at $t=T$ are used as input to the DDIM denoising process [51] to perform editing:
[…]
ControlNet [71] condition $e$ can be added as additional guidance for the sampling and can be obtained from any structured information (e.g., depth maps). The edited frame $\hat{f}^{i}=\mathcal{D}(\hat{z}_{0}^{i})$ is obtained using the decoder $\mathcal{D}$ of the LDM. A classifier-free guidance [24] scale of $s_{cfg}=1$ and a larger scale $s_{cfg}\gg 1$ are used during inversion and denoising, respectively.
B Dataset and Metrics Additional Details
Use Case Number Number of Number of Number of Object Number of Instances Range of Average Instance
of Clips Frames per Clip Objects per Clip Classes per Object Class Mask Size Per Video (%)
MIVE Dataset (full set) 200 $12-46$ 3-12 110 1-20 $0.01\sim 98.68$
For Editing (Things without Stuff) 200 $12-46$ 1-9 54 1-17 $0.02\sim 77.35$
For User Study 30 $15-46$ 4-11 69 1-4 $0.05\sim 80.49$
Table 4: Statistics of our multi-instance video editing dataset in various use cases. (i) MIVE Dataset is the full set of our dataset including both “stuff” and “thing” categories. (ii) For Editing, we only edit objects in the “thing” categories, thus, decreasing some statistics. (iii) For User Study, we only select 30 videos that cover diverse scenarios.
Figure 8: MIVE Dataset caption generation pipeline for each video. Yellow box: The process starts by prompting LLaVA [35] to generate caption for each video that includes all instances in the video. Since LLaVa can only accept images, we perform the prompting for each frame and select one representative caption that includes the most instances. Red box: We utilize Llama 3 [17] to summarize the initial caption generated by LLaVa. Blue box: We manually include all of the instances of interest that are not included in the caption and manually add tags to map the instance captions to corresponding segmentation masks. Purple box: We utilize Llama 3 to generate target captions by retexturing or swapping instances similar to [73] for each instance.
Figure 9: Sample frames and captions from our MIVE Dataset (Part 1). The colored texts are the instance target captions. For each video, the instance masks are color-coded to correspond with the instance target captions in the global target caption.
Figure 10: Sample frames and captions from our MIVE Dataset (Part 2). The colored texts are the instance target captions. For each video, the instance masks are color-coded to correspond with the instance target captions in the global target caption.
B.1 MIVE Dataset Construction
To create our MIVE Dataset, we center-crop a $512\times 512$ region from each video in VIPSeg [39]. We only select the videos where all instances remain visible across frames. We also exclude videos with fewer than 12 frames. To ensure diversity, we remove videos that contain only one of the 40 most frequently occurring object classes (e.g., only persons). This process yields a final subset of 200 videos from VIPSeg.
Since VIPSeg does not include source captions, we generate video captions using a visual language model (LLaVA [35]) and an LLM (Llama 3 [17]). We show our caption generation pipeline in Fig. 8 of this Supplemental. First, we prompt LLaVA to describe the scene of each frame in the video and to include the known objects from the video in the caption. From all the frame captions, we choose one that includes the most instances as the representative caption of the video. We then prompt Llama 3 [17] to summarize this initial caption into a more concise format with fewer tokens and provide the output in JSON format (ommitted in Fig. 8 to simplify visualization). Although LLaVA and Llama provide a useful initial caption for each video, not all objects are accurately captured in each video. Therefore, we manually refine the caption and add the starting and ending tags for each instance token to serve as reference for their corresponding segmentation masks.
To generate a target caption for each instance, we use Llama 3 to prompt edits, such as retexturing or swapping instances similar to [73]. We instruct Llama 3 to generate five candidates for the target caption of each instance and we randomly select one to create a final instance caption. Finally, we modify the original source caption by replacing the source instance captions with the final target instance captions to generate the global target caption. Although we only use the “thing” instances for our task, we still generate captions for the “stuff” objects and background elements. This setup allows for future extensions of our dataset, where an LLM can be employed to create target edits for these objects. We show sample frames and captions in Figures 9 and 10.
B.2 Cross-Instance Accuracy (CIA) Score
Figure 11: Visualization of the Cross-Instance Accuracy (CIA) Score computation. We calculate the Local Textual Faithfulness (LTF) between each cropped instance and all the instance captions. For each row, we assign 1 to the maximum LTF (shown in red) and 0 to the rest. The CIA Score is calculated as the mean of the diagonal elements (shown in green).
Our Cross-Instance Accuracy (CIA) Score is proposed to address the shortcomings of existing video editing metrics, particularly their inability of accounting for potential editing leakage (commonly termed as attention leakage in literatures). The Global Textual Faithfulness (GTF) and Frame Accuracy (FA) metrics cannot account for cases where the text prompt for one instance leaks to another instance, as they cannot capture the nuances of individual instances within the global captions and frames. The Instance Accuracy (IA) only determines whether a cropped instance is more aligned with its target caption compared to its source caption, without considering whether or not another instance caption is affecting the cropped instance. The Local Textual Faithfulness (LTF) only quantifies the alignment of the instance’s target caption and its corresponding cropped instance, but it also overlooks potential attention leakage from other instance captions. While SSIM and LPIPS can measure modifications and attention leakage in the background, they do not take into consideration the leakage in instances that should remain unaffected by a given instance caption. We further observe that the LTF between a cropped instance and another target instance caption, which should not affect the cropped instance, is sometimes higher than the score between the cropped instance and its corresponding target instance caption (red text in Fig. 11).
The nature of our problem, along with the limitations and observations presented above, motivated us to propose a new evaluation metric called the Cross-Instance Accuracy (CIA) Score that can account for attention leakage across instances in video editing tasks. We visualize the computation for our CIA Score in Fig. 11, and provide a detailed explanation of how to calculate the CIA Score in Sec. 4.2 of our main paper.
B.3 Local Metrics Computation
To calculate the local scores, we crop each instance using the bounding boxes inferred from their masks and add padding to preserve their aspect ratios. The Local Textual Faithfulness (LTF) is calculated as the average cosine similarity between the CLIP [43] image embeddings of each cropped instance and the text embeddings of its instance caption, following [55, 1]. Local Temporal Consistency (LTC) is similarly measured as the average cosine similarity between cropped instances across consecutive frames. Instance Accuracy (IA) is the percentage of instances with a higher similarity to the target instance caption than to the source instance caption.
C Comparison with State-of-the-Art Methods
Figure 12: Qualitative comparison for three videos (with increasing difficulty from left to right) in our MIVE dataset. (a) shows the color-coded masks overlaid on the input frames to match the corresponding instance captions. (b)-(f) use global target captions for editing. (g) uses global and instance target captions along with bounding boxes (omitted in (a) for better visualization). (h) uses masks and global and local target captions. Our PRIMEdit in (i) uses instance captions and masks. Unfaithful editing examples are shown in red arrow and attention leakage are shown in green arrow.
C.1 Full Qualitative Comparison and Demo Videos
We present the full qualitative comparison in Fig. 12. Notably, TokenFlow [20] and FreSCo [63] are still susceptible to unfaithful editing as shown in the red arrows in Fig. 12-(f) and (g), respectively.
We provide sample input and edited videos for our method along with baseline methods in our project page: https://kaist-viclab.github.io/primedit-site/.
Method Venue Editing Global Scores Local Scores Leakage Scores User Study
Scope GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ TC $\uparrow$ TF $\uparrow$ Leakage $\uparrow$
ControlVideo [72] ICLR’24 Global 0.9745 0.2738 0.8850 0.9548 0.1960 0.4944 0.4963 0.5327 0.4242 0.0167 0.0656 0.0208
FLATTEN [14] ICLR’24 Global 0.9678 0.2389 0.2627 0.9507 0.1881 0.2463 0.5109 0.8220 0.1314 0.3260 0.0594 0.1646
RAVE [30] CVPR’24 Global 0.9675 0.2728 0.5769 0.9551 0.1869 0.3504 0.4950 0.7699 0.2188 0.0667 0.0354 0.0458
TokenFlow [20] ICLR’24 Global 0.9686 0.2571 0.5617 0.9477 0.1868 0.3499 0.5319 0.7991 0.1807 0.0125 0.0521 0.0531
FreSCo [63] CVPR’24 Global 0.9543 0.2528 0.4202 0.9324 0.1860 0.2960 0.5172 0.7567 0.2427 0.0823 0.0073 0.0104
GAV [28] ICLR’24 Local, Multiple 0.9665 0.2574 0.5544 0.9518 0.1895 0.3733 0.5516 0.8618 0.0914 0.0792 0.0938 0.1000
VideoGrain [65] ICLR’25 Local, Multiple 0.9600 0.2803 0.7430 0.9423 0.2026 0.4798 0.5868 0.8929 0.0488 0.0802 0.1177 0.1542
PRIMEdit (Ours) - Local, Multiple 0.9677 0.2632 0.7707 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586 0.3365 0.5687 0.4510
Table 5:  Full quantitative comparison for multi-instance video editing. The best and second best scores are shown in red and blue, respectively.
C.2 Full Quantitative Results
We present the full quantitative comparison in Tab. 5. Despite achieving the highest scores in local metrics, leakage minimization, background preservation, and user studies, PRIMEdit exhibits a slight disadvantage in global metrics. Our GTC score ranks fourth, which we attribute to prior SOTA methods favoring frame reconstruction, thereby preserving the high GTC of the original video. Similarly, our GTF ranks fourth, while our FA ranks second only to ControlVideo. We attribute this to CLIP’s limitations in compositional reasoning [27, 38], which hinders its ability to capture complex relationships between instances in the global caption and frames. As a result, CLIP-based evaluations may lead to inflated GTF and FA scores despite inaccuracies in instance edits (e.g., swapped edits between two people in Video 2 of Fig. 12-(b)).
C.3 Quantitative Results Based on Instance Sizes and Numbers of Instances
Method Venue Editing Local Scores (Small) Local Scores (Medium) Local Scores (Large)
Scope LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$
ControlVideo [72] ICLR’24 Global 0.9202 0.1610 0.3308 0.9552 0.1993 0.5258 0.9554 0.2008 0.5338
FLATTEN [14] ICLR’24 Global 0.9128 0.1707 0.3228 0.9515 0.1890 0.2366 0.9521 0.1891 0.2285
RAVE [30] CVPR’24 Global 0.9255 0.1710 0.3685 0.9552 0.1898 0.3580 0.9557 0.1898 0.3510
TokenFlow [20] ICLR’24 Global 0.9209 0.1701 0.4015 0.9490 0.1883 0.3451 0.9490 0.1887 0.3405
FreSCo [63] CVPR’24 Global 0.9021 0.1681 0.3686 0.9347 0.1869 0.2948 0.9348 0.1867 0.2825
GAV [28] ICLR’24 Local, Multiple 0.9182 0.1716 0.3358 0.9514 0.1907 0.3810 0.9521 0.1908 0.3807
VideoGrain [65] ICLR’25 Local, Multiple 0.9191 0.1750 0.4419 0.9427 0.2052 0.4991 0.9427 0.2065 0.5025
PRIMEdit (Ours) - Local, Multiple 0.9235 0.1728 0.3870 0.9551 0.2078 0.5502 0.9556 0.2093 0.5604
Table 6:  Quantitative comparison based on instance size. We only show Local Scores since these are the only scores that can be computed depending on the instance size. We follow the categorization of instance size from COCO dataset [34], where: (i) small instance has area $<32^{2}$, (ii) medium instance has area between $32^{2}\text{ and }96^{2}$, and (iii) large instance has area $>96^{2}$. The best and second best scores are shown in red and blue, respectively.
Method Venue Editing Global Scores Local Scores Leakage Scores
Scope GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Editing on One (1) Instance (Easy Video) - 200 Videos
ControlVideo [72] ICLR’24 Global 0.9755 0.2717 0.6922 0.9611 0.2048 0.5351 1.0000 0.4130 0.5090
FLATTEN [14] ICLR’24 Global 0.9699 0.2538 0.3140 0.9586 0.1826 0.1971 1.0000 0.7783 0.1567
RAVE [30] CVPR’24 Global 0.9691 0.2774 0.4243 0.9599 0.1857 0.3374 1.0000 0.7044 0.2716
TokenFlow [20] ICLR’24 Global 0.9715 0.2621 0.4413 0.9564 0.1822 0.3006 1.0000 0.7546 0.2148
FreSCo [63] CVPR’24 Global 0.9563 0.2630 0.3200 0.9423 0.1805 0.2515 1.0000 0.6894 0.3014
GAV [28] ICLR’24 Local, Multiple 0.9701 0.2537 0.4453 0.9584 0.1851 0.3350 1.0000 0.8340 0.0999
VideoGrain [65] ICLR’25 Local, Multiple 0.9659 0.2696 0.5474 0.9515 0.2020 0.4531 1.0000 0.8731 0.0570
PRIMEdit (Ours) - Local, Multiple 0.9737 0.2534 0.6192 0.9614 0.2108 0.5562 1.0000 0.8807 0.0661
Editing on 1-3 Instances (Medium Video) - 116 Videos
ControlVideo [72] ICLR’24 Global 0.9730 0.2724 0.8839 0.9512 0.2020 0.5380 0.6191 0.4879 0.4515
FLATTEN [14] ICLR’24 Global 0.9659 0.2417 0.2908 0.9484 0.1894 0.2557 0.6053 0.7902 0.1504
RAVE [30] CVPR’24 Global 0.9661 0.2699 0.5437 0.9532 0.1886 0.3604 0.5971 0.7396 0.2379
TokenFlow [20] ICLR’24 Global 0.9686 0.2580 0.5701 0.9463 0.1880 0.3476 0.6265 0.7720 0.1974
FreSCo [63] CVPR’24 Global 0.9535 0.2492 0.4140 0.9326 0.1877 0.2836 0.6084 0.7165 0.2811
GAV [28] ICLR’24 Local, Multiple 0.9649 0.2528 0.5648 0.9481 0.1919 0.3903 0.6498 0.8397 0.0959
VideoGrain [65] ICLR’25 Local, Multiple 0.9596 0.2775 0.7486 0.9413 0.2065 0.5238 0.6712 0.8736 0.0526
PRIMEdit (Ours) - Local, Multiple 0.9658 0.2629 0.7897 0.9516 0.2096 0.5817 0.7567 0.8823 0.0633
Editing on 4-7 Instances (Hard Video) - 66 Videos
ControlVideo [72] ICLR’24 Global 0.9760 0.2774 0.8815 0.9581 0.1874 0.4544 0.3694 0.5830 0.3920
FLATTEN [14] ICLR’24 Global 0.9702 0.2372 0.2303 0.9527 0.1854 0.2288 0.4206 0.8585 0.1099
RAVE [30] CVPR’24 Global 0.9690 0.2776 0.5821 0.9571 0.1843 0.3399 0.3954 0.8060 0.1957
TokenFlow [20] ICLR’24 Global 0.9686 0.2559 0.5427 0.9487 0.1845 0.3631 0.4504 0.8342 0.1585
FreSCo [63] CVPR’24 Global 0.9557 0.2589 0.4274 0.9319 0.1835 0.3179 0.4391 0.8068 0.1934
GAV [28] ICLR’24 Local, Multiple 0.9679 0.2652 0.5557 0.9553 0.1861 0.3659 0.4692 0.8884 0.0837
VideoGrain [65] ICLR’25 Local, Multiple 0.9603 0.2836 0.6938 0.9429 0.1972 0.4185 0.5096 0.9180 0.0425
PRIMEdit (Ours) - Local, Multiple 0.9691 0.2622 0.7039 0.9578 0.1990 0.4897 0.6087 0.9229 0.0509
Editing on $>$7 Instances (Extreme Video) - 18 Videos
ControlVideo [72] ICLR’24 Global 0.9781 0.2692 0.9051 0.9651 0.1885 0.3602 0.1698 0.6374 0.3660
FLATTEN [14] ICLR’24 Global 0.9717 0.2274 0.2007 0.9584 0.1898 0.2499 0.2339 0.8930 0.0881
RAVE [30] CVPR’24 Global 0.9711 0.2735 0.7714 0.9602 0.1857 0.3242 0.2026 0.8323 0.1806
TokenFlow [20] ICLR’24 Global 0.9682 0.2552 0.5778 0.9529 0.1878 0.3162 0.2212 0.8455 0.1545
FreSCo [63] CVPR’24 Global 0.9538 0.2536 0.4339 0.9322 0.1844 0.2954 0.2156 0.8319 0.1760
GAV [28] ICLR’24 Local, Multiple 0.9715 0.2580 0.4825 0.9628 0.1870 0.2911 0.2210 0.9070 0.0912
VideoGrain [65] ICLR’25 Local, Multiple 0.9615 0.2873 0.9286 0.9478 0.1966 0.4050 0.2516 0.9348 0.0463
PRIMEdit (Ours) - Local, Multiple 0.9750 0.2682 0.8937 0.9689 0.1945 0.4209 0.3413 0.9390 0.0559
Table 7:  Quantitative comparison for multi-instance video editing on various number of instances. We categorize 200 videos of MIVE Dataset depending on the number of edited instances: (i) Easy Video (EV): video that contains 1 edited instance, (ii) Medium Video (MV): video that contains 1-3 edited instances, (iii) Hard Video (HV): video that contains 4-7 edited instances, and (iv) Extreme Video (XV): video that contains $>$ 7 edited instances. The best and second best scores are shown in red and blue, respectively.
In this part, we analyze whether the baseline methods and our PRIMEdit exhibit biases based on: (i) the instance sizes or (ii) the number of instances.
Quantitative results based on instance sizes. Table 6 presents a comparison between our method and baselines across various instance sizes. To categorize instance sizes, we follow the COCO dataset [34], where: (i) small instances have the areas $<32^{2}$, (ii) medium instances have the areas between $32^{2}$ and $96^{2}$, and (iii) large instances have the areas $>96^{2}$. Across all videos, there are 69 small instances, 297 medium instances, and 434 large instances. To compute the area of each instance in the video, we average its area across all frames.
As shown in Tab. 6, the Local Temporal Consistency scores closely align with the findings in the main paper (Sec. 5.1).
For medium and large instance sizes, our method achieves the highest performance in textual faithfulness (LTF and IA) and the second-best in temporal consistency. In these scenarios, both our PRIMEdit and VideoGrain show significant improvements over other SOTA methods in terms of LTF and IA. Notably, in certain cases, global video editing methods outperform GAV. These results underscore the advantage of using masks over bounding boxes as conditioning mechanisms for multi-instance video editing, enabling more precise and context-aware modifications.
In the small instance size scenario, the LTF scores of all methods tend to be lower compared to the medium and large instances. This highlights the challenges of editing small instances in diffusion-based video editing methods, caused by the downsampling in the VAE of the LDM [46].
Quantitative results based on number of instances. In addition to the quantitative comparisons with respect to various instance sizes, we provide further analysis for the different numbers of edited instances in each video, as shown in Tab. 7. Videos are categorized into four groups: (i) easy video (EV) having only one edited instance, (ii) medium video (MV) having 1–3 edited instances, (iii) hard video (HV) having 3-4 edited instances, and (iv) extreme video (XV) having 7 or more edited instances.
For the Leakage Scores, there is no significant deviation from the quantitative comparisons in the main paper (Tab. 2). In most cases, our PRIMEdit achieves the best Leakage Scores performance, while VideoGrain always secures the second-best performance. For Local Scores, the results are consistent with the main paper: our PRIMEdit almost always achieves the best LTC, LTF, and IA scores. Similarly, the global scores align with Tab. 5, where our method achieves competitive GTC and secures the second-best FA performance.
Notably, our PRIMEdit’s performance gain is not solely due to multi-instance editing. PRIMEdit also excels in single-instance editing, as evidenced in the results of EV, where we ourperform baselines specifically optimized for single-object editing across all local metrics. This underscores PRIMEdit’s versatility and effectiveness in delivering localized and faithful edits across various scenarios.
In summary, while there are minor deviations in some scores compared to the quantitative results in Tab. 5 and in Tab. 7, the performance of our PRIMEdit remains consistent across various scenarios. This demonstrates the robustness of our approach against variations in instance sizes and the number of edited instances.
C.4 User Study Details
Figure 13: Our user study interface and questionnaire form. Participants are presented with an input video with a source caption, an annotated video with a target caption, and 8 randomly ordered videos edited using our PRIMEdit and seven other SOTA video editing methods. Each instance mask in the annotated video is color-coded to correspond with its instance target caption. Participants are tasked to select the video with the Best Temporal Consistency, Best Textual Faithfulness, and Least Editing Leakage.
To conduct our user study, we select 30 videos from our dataset covering diverse scenarios (varying numbers of objects per clip, numbers of instances per object class, and instance sizes). We show the statistics of the videos we used in our user study in Tab. 4. We edit the videos using our PRIMEdit framework and the other seven SOTA video editing methods, namely, ControlVideo [72], FLATTEN [14], RAVE [30], TokenFlow [20], FreSCo [63], GAV [28], and VideoGrain [65]. We task 32 participants to select the method with the:
  • Best Temporal Consistency: Choose the video with the smoothest transitions;
  • Best Textual Faithfulness: Choose the video with the most accurate text-object alignment. Make sure to check the video’s alignment with the overall target caption and the individual instance captions;
  • Least Editing Leakage: Choose the video with the least text leakage onto other objects and/or the background.
Before starting the user study, we provided the participants with good and bad examples for each criterion for guidance. Fig. 13 shows our user study interface and the questionnaire form.
C.5 Attention Visualization
Figure 14: Attention weight visualization.
In Fig. 14, we present the attention weights for the editing results shown in Fig. 6 of the main paper. Our IPR reduces artifacts better than DenseDiffusion [32] by smoothing attention from the start (S) to the end of the text token (E), whereas DenseDiffusion concentrates on the text token (T), making it more prone to artifacts (Fig. 6-(c) of the main paper).
C.6 Video-P2P Results
Figure 15: Video-P2P [36] results on recursive multi-instance editing. The artifacts that accumulate when Video-P2P is used repeatedly for multi-instance editing is shown in red arrow. Our PRIMEdit prevents this error accumulation since we do not edit the frames recursively.
While the recent work Video-P2P [36] is capable of local video editing, its design limits it to editing only one object at a time. Attempting to edit multiple objects recursively with this method results in artifacts, as demonstrated in Fig. 15. Consequently, we exclude Video-P2P from our SOTA comparison.
D Additional Analysis and Ablation Studies
Methods GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
IPR No Modulation [46] 0.9680 0.2614 0.7324 0.9564 0.2018 0.4920 0.6497 0.9007 0.0600
Dense Diffusion [32] 0.9699 0.2646 0.8272 0.9530 0.2078 0.5845 0.6774 0.8977 0.0604
Ours, Full 0.9677 0.2632 0.7707 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
DMS Only SNS 0.9641 0.2623 0.7661 0.9503 0.2047 0.5288 0.6743 0.9063 0.0500
Only PNS 0.9670 0.2638 0.8008 0.9542 0.2044 0.5311 0.6718 0.8993 0.0585
SNS + PNS (no re-inv) 0.9665 0.2624 0.7688 0.9535 0.2051 0.5323 0.6728 0.9040 0.0525
Ours, Full 0.9677 0.2632 0.7707 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
Table 8: The full results of our ablation study on Disentangled Multi-instance Sampling (DMS) and Instance-centric Probability Redistribution (IPR). LPS and NPS denotes Latent Parallel Sampling and Noise Parallel Sampling, respectively. The best and second best scores are shown in red and blue, respectively.
We present the Global Scores of in Tab. 8, which could not be included in the ablation study of the main paper due to space limitations.
In the ablation study on IPR, the Global Scores align with the Local Scores, with DenseDiffusion achieving better global editing faithfulness performance. However, as shown in Fig. 6 of the main paper, Dense Diffusion may exhibit artifacts. These artifacts reduce the local scores but do not impact the global scores. In some cases, they may even improve the global scores, as global scores evaluate each frame with the whole word tokens of the caption. If each token is visible in any position of the frame, it can lead to higher global scores due to CLIP’s limitations in compositional reasoning [27, 38].
In the ablation study on DMS, our full method achieves the highest GTC and competitive global editing faithfulness (GTF and FA), while excelling in local metrics with a slight reduction in background preservation scores. This demonstrates the effectiveness of our approach in balancing global and local coherence while maintaining strong background preservation.
D.1 Analysis on IPR
Figure 16: IPR analysis: Effect of altering the attention probability values of padding token $A_{i,j\in\textbf{P}}$.
Figure 17: IPR analysis: Various scenarios of redistributing the attention probability values of tokens $S$, T, and $E$. Row 2-4: Redistributing the probability values from either T or $E$ tokens to $S$ decreases the editing faithfulness. Row 5-6: Redistributing the probability values from $S$ token by a constant factor $\lambda_{S}$ to the T and $E$ tokens improves the editing faithfulness. Ours: Redistributing the probability value of the $S$ token using our proposed dynamic approach achieves the best editing faithfulness.
Our Instance-centric Probability Redistribution (IPR) is proposed based on two key observations: (i) manipulating the attention probabilities of the padding tokens $A_{i,j\in\boldsymbol{P}}$ may lead to artifacts; therefore, we avoid altering them; and (ii) increasing the $S$ token’s probability decreases editing faithfulness, whereas decreasing it and reallocating those values to the $\boldsymbol{T}$ and $E$ tokens improves editing faithfulness. Figure 16 illustrates the first observation, showing that altering the padding tokens tends to cause artifacts, e.g., the alien face exhibiting noise and oversaturation. In contrast, our proposed IPR avoids modifying the padding tokens $A_{i,j\in\boldsymbol{P}}$, preventing oversaturation and producing cleaner editing results.
The second observation concerns the redistribution of attention probability values among the $S$, $\boldsymbol{T}$, and $E$ tokens. Figure 17 illustrates the results of different redistribution approaches. We observe the following: (i) Redistributing attention probability values from both the $\boldsymbol{T}$ and $E$ tokens to $S$ (second row), as well as from the $\boldsymbol{T}$ token to $S$ (third row), reduces editing faithfulness. For example, the right washing machine appears gray and does not transform into a yellow washing machine, and the person remains human-like instead of transforming into an alien. Moreover, while the structure of the left washing machine is starting to change, it still does not resemble an oven. (ii) Redistributing attention probability values from the $E$ token to $S$ (fourth row) causes the right washing machine to transform into a yellow washing machine, but neither the left washing machine nor the person transforms into an oven or an alien, respectively. (iii) Redistributing attention probability values from the $S$ token to both the $\boldsymbol{T}$ and $E$ tokens by setting $\lambda_{S}=0.2$ (fifth row) begins to achieve more faithful editing. For instance, the human’s eyes start to resemble alien eyes, the left washing machine begins to transform into an oven, and the right washing machine starts to turn into a yellow washing machine. These first three observations highlight the importance of redistributing attention probabilities from $S$ to both $\boldsymbol{T}$ and $E$ tokens, which forms the main motivation for our proposed IPR. (iv) Both increasing $\lambda_{S}$ and utilizing our IPR to compute $\lambda_{S}$ enhance editing faithfulness. Notably, our proposed IPR with dynamic $\lambda_{S}$ achieves better editing faithfulness, e.g., improved faithfulness in both the alien’s face, oven, and yellow washing machine.
D.2 IPR Ablations
Method Global Scores Local Scores Leakage Scores
GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
(a) Ablation on Percentage of Sampling Step that Utilizes IPR
Applying IPR on The First $10\%$ Sampling Steps (Ours) 0.9660 0.2636 0.7773 0.9528 0.2060 0.5390 0.6806 0.9055 0.0503
Applying IPR on The First $20\%$ Sampling Steps 0.9657 0.2645 0.7825 0.9524 0.2062 0.5387 0.6783 0.9053 0.0502
Applying IPR on The First $30\%$ Sampling Steps 0.9656 0.2648 0.7923 0.9516 0.2062 0.5494 0.6782 0.9052 0.0501
(b) Ablation on $\lambda$
$\lambda=0.4$ 0.9658 0.2621 0.7635 0.9532 0.2048 0.5345 0.6793 0.9053 0.0506
$\lambda=0.5$ (Ours) 0.9658 0.2623 0.7668 0.9531 0.2050 0.5365 0.6823 0.9054 0.0505
$\lambda=0.6$ 0.9658 0.2627 0.7685 0.9531 0.2053 0.5334 0.6791 0.9054 0.0504
$\lambda=0.7$ 0.9659 0.2632 0.7732 0.9530 0.2057 0.5363 0.6807 0.9055 0.0503
(c) Ablation on $\lambda_{r}$
$\lambda_{r}=0.4$ 0.9659 0.2635 0.7745 0.9530 0.2053 0.5360 0.6764 0.9052 0.0513
$\lambda_{r}=0.5$ (Ours) 0.9659 0.2625 0.7692 0.9532 0.2051 0.5349 0.6816 0.9053 0.0509
$\lambda_{r}=0.6$ 0.9658 0.2623 0.7668 0.9531 0.2050 0.5365 0.6823 0.9054 0.0505
$\lambda_{r}=0.7$ 0.9660 0.2625 0.7737 0.9530 0.2053 0.5375 0.6807 0.9055 0.0502
Table 9:  Ablation study on various hyperparameter configurations for our Instance-centric Probability Redistribution (IPR). The best and second best scores are shown in red and blue, respectively.
Figure 18: Ablation study on IPR: (a) percentage of sampling steps where we apply our IPR. Increasing the IPR step percentage generally enhances editing faithfulness; however, applying it excessively can lead to the loss of object structures (e.g., the alien’s face and bag strap). Our results indicate that applying IPR for 10% of the sampling steps provides the best trade-off between editing faithfulness and structural preservation.
Figure 19: Ablation study on IPR: (b) $\lambda$. Increasing $\lambda$ steers the editing more towards the edit caption, but increasing it too much to $\lambda=0.7$ may cause oversaturation, e.g., the alien’s face and clothes. The best trade-off is $\lambda=0.5$.
Figure 20: Ablation study on IPR: (c) $\lambda_{r}$. Increasing $\lambda_{r}$ enhances details (e.g., the alien’s face) but the object tends to revert back to the original object. The best trade-off between editing faithfulness and enhancement in details is $\lambda_{r}=0.5$.
In this subsection, we perform additional ablation studies on our IPR, focusing on the hyperparameters $\lambda$, $\lambda_{r}$, and the percentage of DDIM steps where IPR is applied. The quantitative comparison is presented in Tab. 9 and qualitative comparisons are presented in Fig. 18, Fig. 19, and Fig. 20.
The first hyperparameter we determine is the percentage of steps during which we apply our IPR (IPR steps percentage). Figure 18 illustrates the qualitative results of different configurations for the IPR steps percentage. Increasing the IPR steps percentage generally enhances editing faithfulness, e.g., better facial texture details in the alien’s head. However, excessively increasing the percentage may lead to loss of structure. Our quantitative results in Tab. 9 align with these findings, showing that increasing the IPR steps percentage tends to improve editing faithfulness metrics (GTF, LTF, FA, and IA). Based on these observations, we choose to apply IPR to 10% of the sampling steps.
Figure 19 illustrates the results of varying the values of the hyperparameter $\lambda$. From these results, we observe that increasing $\lambda$ steers the editing more towards the edit caption. For example, the alien’s face is slowly losing the man’s features as we increase $\lambda$. However, excessively increasing $\lambda$ may cause oversaturation as observed in the alien’s face and clothes. In terms of quantitative performance, both $\lambda=0.5$ and $\lambda=0.6$ yield comparable results. However, due to the potential risk of oversaturation, we use $\lambda=0.5$ as our default.
For the next hyperparameter, $\lambda_{r}$, in our IPR, the qualitative results are shown in Fig. 20. We observe that increasing $\lambda_{r}$ tends to increase details but introduces artifacts. For example,while the alien’s face becomes more detailed, it also exhibits noticeable noise, partially reverting to its original human form. These observations are further supported by the quantitative comparison in Tab. 9, where increasing $\lambda_{r}$ reduces faithfulness scores (GTF, FA, LTF, and IA). Based on these findings, we select $\lambda_{r}=0.5$ to achieve a balance between overall editing faithfulness and the minimization of artifacts.
Overall, the selection of these three hyperparameters in IPR is guided by the trade-off between editing faithfulness, emergence of artifacts, and oversaturation.
D.3 DMS Ablations
Figure 21: Ablation study on DMS: (a) Ablation on the number of SNS steps. Increasing the number of SNS steps reduces artifacts in the background and improves local faithfulness and background preservation scores. Setting SNS steps to the first 40% of the sampling process achieves a balanced trade-off between qualitative and quantitative performance. See Tab. 10-(a) for quantitative results.
Figure 22: Ablation study on DMS: (b) Ablation on the number of re-inversion steps $l$. Increasing the number of re-inversion steps enhances the global scores as well as the local temporal consistency performance of PRIMEdit. Quantitative results are provided in Tab. 10-(b). However, we set the number of re-inversion steps to $l=2$ to avoid blurring artifacts shown in (4).
Method Global Scores Local Scores Leakage Scores
GTC $\uparrow$ GTF $\uparrow$ FA $\uparrow$ LTC $\uparrow$ LTF $\uparrow$ IA $\uparrow$ CIA (Ours) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
(a) Ablation on the number of SNS steps
(1) No SNS (Pure PNS) 0.9670 0.2638 0.8008 0.9542 0.2044 0.5311 0.6718 0.8993 0.0585
(2) First 10% 0.9667 0.2634 0.7855 0.9541 0.2047 0.5356 0.6704 0.9009 0.0565
(3) First 20% 0.9663 0.2635 0.7780 0.9535 0.2050 0.5409 0.6739 0.9021 0.0551
(4) First 30% 0.9662 0.2631 0.7687 0.9534 0.2052 0.5390 0.6718 0.9031 0.0537
(5) First 40% (Ours) 0.9665 0.2624 0.7688 0.9535 0.2051 0.5323 0.6728 0.9040 0.0525
(b) Ablation on the number of re-inversion steps
(1) l = 0 0.9665 0.2624 0.7688 0.9535 0.2051 0.5323 0.6728 0.9040 0.0525
(2) l = 1 0.9671 0.2629 0.7710 0.9544 0.2050 0.5319 0.6713 0.9029 0.0549
(3) l = 2 (Ours) 0.9677 0.2632 0.7707 0.9552 0.2048 0.5369 0.6705 0.9008 0.0586
(4) l = 3 0.9682 0.2635 0.7734 0.9560 0.2045 0.5342 0.6714 0.8976 0.0637
Table 10: Ablation study on various hyperparameter configurations for our Disentangled Multi-instance Sampling (DMS). The best and second best scores are shown in red and blue, respectively.
We conduct ablation studies on the hyperparameters of our Disentangled Multi-instance Sampling (DMS) technique. The quantitative results are summarized in Tab. 10, while the qualitative comparisons are shown in Figures 21 and 22.
First, we determine the optimal number of SNS steps during sampling. Without SNS, the framework struggles with background preservation and introduces artifacts (Fig. 21-(1)). As the number of SNS steps increases, artifacts gradually decrease (Fig. 21-(2) to (5)), leading to improved local faithfulness and background preservation scores, though with a slight decrease in temporal consistency. Based on our findings, setting SNS steps to the first 40% of the sampling process achieves a balanced trade-off between qualitative and quantitative performance (Tab. 10-(a)).
Finally, we conduct experiments by introducing re-inversion between the SNS and PNS steps. Increasing the number of re-inversion steps $l$ after SNS improves global metrics as well as the local temporal consistency (Tab. 10-(b)). However, excessively increasing $l$ leads to blurring and artifacts Fig. 22-(4). Ultimately, we set $l=2$ to balance the overall qualitative appearance and quantitative performance.
E Qualitative Results on DAVIS Dataset
Figure 23: Qualitative results on four DAVIS Dataset videos. Our PRIMEdit maintains faithful and temporally consistent edits, particularly in cases with large object motions (rows 3 and 4).
Fig. 23 presents qualitative results on four videos from the DAVIS dataset, demonstrating the effectiveness of our PRIMEdit in maintaining faithful and temporally consistent edits. Our approach preserves object details, even in challenging scenarios involving large object motions (rows 3 and 4). These results highlight the robustness of our method in achieving high-quality video edits with strong temporal coherence.
F Efficiency Comparison with VideoGrain
Method Time (sec) $\downarrow$ Memory (GB) $\downarrow$
Single Instance, 15 frames
VideoGrain [65] 328.28 21.96
PRIMEdit (Ours) 143.27 15.28
Single Instance, 32 frames
VideoGrain [65] 1181.41 41.90
PRIMEdit (Ours) 295.43 24.24
3 Instances, 15 frames
VideoGrain [65] 331.39 22.86
PRIMEdit (Ours) 225.71 15.33
3 Instances, 32 frames
VideoGrain [65] 1181.08 46.06
PRIMEdit (Ours) 465.06 24.3
Table 11: Runtime and GPU memory usage comparison between our PRIMEdit and VideoGrain [65]. The best and second best scores are shown in red and blue, respectively.
We compare runtime and GPU memory usage against VideoGrain [65] in Tab. 11, using a single RTX A6000 GPU for editing under four scenarios: (i) a single instance over 15 frames, (ii) a single instance over 32 frames, (iii) three instances over 15 frames, and (iv) three instances over 32 frames. As shown in Tab. 11, our PRIMEdit framework is both faster and more memory-efficient for multi-instance video editing.
G Limitations
Figure 24: Limitation. Our method struggles in scenes with reflective surfaces.
As shown in Fig. 24, our model struggles with reflection consistency due to mask limitations, covering only the instance, not its reflection. This causes discrepancies between the edited instance and its reflection in scenes with reflective surfaces. Future works can build on approaches like [16] to ensure consistency for both the instance and its reflection.